OpenStudy (anonymous):

Use the equation and given derivative to find the specified derivative. y=3x+5 given that dx/dt=2, find dy/dt when x=1

OpenStudy (anonymous):

\[y=3x+5\] \[\frac{dy}{dt}=\frac{dy}{dx}\frac{dx}{dt}\]

OpenStudy (anonymous):

you know \[\frac{dy}{dx}=3\] and you are also told that \[\frac{dx}{dt}=2\] and this makes \[\frac{dy}{dt}=3\times 2=6\] a constant, so it is 6 no matter what x is

OpenStudy (anonymous):

thank you!